OpenStudy (anonymous):

one number is 7 more than twice the other number.If the sum of the two numbers is 37, find the numbers ! Please help in need !!

jimthompson5910 (jim_thompson5910):

let x and y be the two numbers

jimthompson5910 (jim_thompson5910):

"one number is 7 more than twice the other number" so that means y = 2x+7

jimthompson5910 (jim_thompson5910):

"the sum of the two numbers is 37" meaning that x+y = 37

OpenStudy (anonymous):

so the first equation is 2x+7 ??

jimthompson5910 (jim_thompson5910):

y = 2x+7 you need the equal sign to have it be an equation

jimthompson5910 (jim_thompson5910):

x+y = 37 x+2x+7 = 37 ... replace y with 2x+7 (works because y = 2x+7) 3x+7 = 37 x = ???

OpenStudy (anonymous):

i have 2x^2 +7x-37=0 but that doesnt factor

OpenStudy (anonymous):

x=3

jimthompson5910 (jim_thompson5910):

that's incorrect

jimthompson5910 (jim_thompson5910):

x = 3 is correct though

jimthompson5910 (jim_thompson5910):

3x+7 = 37 leads to x = 3

jimthompson5910 (jim_thompson5910):

if x = 3, then y = 2x+7 = ???

OpenStudy (anonymous):

so the other # is 13?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

thank you!

jimthompson5910 (jim_thompson5910):

actually hold on

jimthompson5910 (jim_thompson5910):

sorry I jumped the gun

jimthompson5910 (jim_thompson5910):

3x+7 = 37 leads to x = 10 and NOT x = 3

jimthompson5910 (jim_thompson5910):

you should have 3x = 30 ---> x = 10

jimthompson5910 (jim_thompson5910):

if x = 10, then y = 2x+7 = 2(10)+7 = 27

OpenStudy (anonymous):

okay so 10 and 27

OpenStudy (anonymous):

thank you !

jimthompson5910 (jim_thompson5910):

yes and you check by adding them: 10+27 = 37